A tungsten wire has a radius of 0.078 mm and is heated from 20.0 to 1369 oC. The temperature coefficient of resistivity is α = 4.5 x 10-3 (Co)-1. When 140 V is applied across the ends of the hot wire, a current of 2.2 A is produced. How long is the wire? Neglect any effects due to thermal expansion of the wire.
Answer #1

Ohm's law applies

R=\frac{V}{I}=\frac{ 140V}{2.2A }=63.64\Omega hot resistance

New resistivity is calculated

\rho =\rho _{0}\left [ 1\dotplus \alpha \left ( T_{f}-T_{i} \right ) \right ]

where:

resistivity at 20 degrees Celsius

\rho _{0}=5.6\cdot 10^{-8}\Omega \cdot m

initial temperature

t_{i}=20^{0}C
final temperature

t_{f}=1369^{0}C
The temperature coefficient of resistivity

\alpha =4.5\cdot 10^{-3}^0C^{-1}

evaluate numerically

\rho =\rho _{0}\left [ 1\dotplus \alpha \left ( T_{f}-T_{i} \right ) \right ]=5.6\cdot 10^{-8}\Omega \cdot m\cdot \left [ 1\dotplus 4.5\cdot 10^{-3}..^0C^{-1}\cdot \left (1369^{0}C-20.0^{0}C \right )\right ]=4.0\cdot 10^{-7}\Omega \cdot m

\rho =4.0\cdot 10^{-7}\Omega \cdot m

The resistance is given by

R=\rho \cdot \frac{L}{A}

where:
hot resistance

R=63.64\Omega
hot resistivity

\rho =4.0\cdot 10^{-7}\Omega \cdot m
long

L
area

A=\pi \cdot r^{2}=\pi \cdot \left ( 7.8\cdot 10^{-8}m \right )^{2}=1.9\cdot 10^{-8}m^{2}

r=0.078mm\cdot \frac{1m }{1000mm }=7.8\cdot 10^{-8}m

expression for L

R=\rho \cdot \frac{L}{A}\Rightarrow L=\frac{R\cdot A}{\rho }

evaluate numerically

L=\frac{R\cdot A}{\rho }=\frac{63.64\Omega\cdot 1.9\cdot 10^{-8}m^{2}}{4.0\cdot 10^{-7}\Omega \cdot m }=3.0m